OpenStudy (anonymous):

Need help with a problem. Just tell me where to go with it; I don't want anyone to solve it for me. Just confused... I tried the counting principle but it didn't work. "Suppose you roll two number cubes and pick a letter of the alphabet at random. Find the probability you roll two even numbers and pick one of the vowels a, e, i, o, or u"

OpenStudy (anonymous):

what is the probability of rolling an even number on one of the cubes? What is the probability of choosing one of the vowels? I mean on there own

OpenStudy (anonymous):

Probability of choosing an even number on either cube is 1/2, and the probability for choosing a vowel is 5/26. Do I just add them together?

OpenStudy (anonymous):

are they independent events?

OpenStudy (anonymous):

...Yes, I think.

OpenStudy (anonymous):

well they cannot impact each other so they are independent events

OpenStudy (anonymous):

right. And the probability for rolling to even numbers is 1/4

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

Unfortunately I do not remeber the formula for independent events so I would check you textbook or formula sheet . . . if I had to guess I would say it is multiplying them together so \[\frac{ 1 }{ 2 }*\frac{ 1 }{ 2 }*\frac{ 5 }{ 26 }\]

OpenStudy (anonymous):

Let's see...

OpenStudy (anonymous):

Ah! That's it. Thanks.

OpenStudy (anonymous):

yea a quick way to see that it is not adding is that it would give you over 100% chance of success anyways don't forget to close the question